OpenStudy (anonymous):

ok my problem is 1 5/8 + 1 3/4 = 3 3/8 How?

OpenStudy (anonymous):

1 5 / 8 is 13/8 1 3 /4 is 7 / 4 which is 14 /8 13/8 + 14 /8 = 27/8 3 3/8 is 27/8

OpenStudy (anonymous):

Here we have to find the least common multiple to add both the term so, 1*5/8=8*1+5/8=13/8 similarly 1*3/4=1*4+3/4=7/4 now, 13/8+7/4=13+14/8=27/8 and we can write it like....3*3/8=3*8+3/8

OpenStudy (anonymous):

hey do you know david

OpenStudy (anonymous):

yassou

OpenStudy (anonymous):

thank you both!